Consider the following primal problem;


minimize x0=20x1+30x2+50x3+40x4

subject to:

4x1+6x2+x3+2x4≥12

2x1+x2+6x3+5x4≥14

x1+2x2+4x3+3x4≥8

xi≥0,i=1,2,3,4

Use simplex method to solve the dual of the above primal problem


1
Expert's answer
2021-10-18T12:36:31-0400

"MIN \\ Z_x\t=\t\t20\tx_1\t+\t30\tx_2\t+\t50\tx_3\t+\t40\tx_4\\\\\nSubject \\ to:\\\\\n4\tx_1\t+\t6\tx_2\t+\t\tx_3\t+\t2\tx_4\t\u2265\t12\\\\\n2\tx_1\t+\t\tx_2\t+\t6\tx_3\t+\t5\tx_4\t\u2265\t14\\\\\nx_1\t+\t2\tx_2\t+\t4\tx_3\t+\t3\tx_4\t\u2265\t8\\\\\nand \\ x_1,x_2,x_3,x_4\u22650;"

"Dual \\ is \\ (Solution \\ steps \\ of \\ Dual\\ by\\ BigM \\ method)\\\\\n\nMAX\\ Z_y\t=\t\t12\ty_1\t+\t14\ty_2\t+\t8\ty_3\\\\\nSubject\\ to:\\\\\n4\ty_1\t+\t2\ty_2\t+\t\ty_3\t\u2264\t20\\\\\n6\ty_1\t+\t\ty_2\t+\t2\ty_3\t\u2264\t30\\\\\ny_1\t+\t6\ty_2\t+\t4\ty_3\t\u2264\t50\\\\\n2\ty_1\t+\t5\ty_2\t+\t3\ty_3\t\u2264\t40\\\\\nand\\ y_1,y_2,y_3\u22650;\\\\"

"Max: Z=12 y_{1}+14 y_{2}+8 y_{3} \\subject \ to: 4 y_{1}+2 y_{2}+y_{3} \leq 20 \\ 6 y_{1}+y_{2}+2 y_{3} \leq 30 \\y_{1}+6 y_{2}+4 y_{3} \leq 50 \\ 2 y_{1}+5 y_{2}+3 y_{3} \leq 40 \\ and \ y_{1}, y_{2}, y_{3} \geq 0 \\"

The problem is converted to canonical form by adding slack, surplus, and artificial variables as appropriate


1. As the constraint-1 is of type '≤' we should add slack variable "S_1"


2. As the constraint-2 is of type '≤' we should add slack variable "S_2"


3. As the constraint-3 is of type '≤' we should add slack variable "S_3"


4. As the constraint-4 is of type '≤' we should add slack variable "S_4"


After introducing slack variables


"\operatorname{Max} Z=12 y_{1}+14 y_{2}+8 y_{3}+0 S_{1}+0 S_{2}+0 S_{3}+0 S_{4} \\
subject \ to:\\
 
\begin{aligned}
4 y_{1}+2 y_{2}+y_{3}+S_{1} =20 \\
6 y_{1}+y_{2}+2 y_{3}+S_{2} =30 \\
y_{1}+6 y_{2}+4 y_{3} +S_3=50 \\
2 y_{1}+5 y_{2}+3 y_{3} +S_{4} =40\\
\end{aligned}\\
 
 
\text { and } y_{1}, y_{2}, y_{3}, S_{1}, S_{2}, S_{3}, S_{4} \geq 0"




The negative minimum Zj-Cj is -14 and its column index is 2. So, the entering variable is "y_2" .


The minimum ratio is 8 and its row index is 4. So, the leaving basis variable is "S_4" .


∴ The pivot element is 5.


Entering ="y_2" , Departing ="S_4" , Key Element =5





The negative minimum Zj-Cj is -6.4 and its column index is 1. So, the entering variable is y1.


The minimum ratio is 1.25 and its row index is 1. So, the leaving basis variable is S1.


 ∴ The pivot element is 3.2.


Entering =y1, Departing =S1, Key Element =3.2






Since all Zj-Cj≥0


Hence, the optimal solution arrives with value of variables as :

"y_1=1.25,y_2=7.5,y_3=0"


Max Z=120